OpenStudy (anonymous):

Solve for t 7^(t+1)=8^t

OpenStudy (turingtest):

take the log of both sides and remember that\[\log(a^b)=b\log a\]

OpenStudy (anonymous):

I actually don't know what that means...

OpenStudy (turingtest):

it doesn't matter what base the log is

OpenStudy (anonymous):

you have to intro duce lgas both sides

OpenStudy (turingtest):

@megachomehead6 do you not know what a logarithm is?

OpenStudy (anonymous):

No, my professor threw this section on us without teaching it to us

OpenStudy (anonymous):

you are in college and you never heard of logarithm,what about ln

OpenStudy (anonymous):

Sorry, yes I've heard of logs but not in this form

OpenStudy (turingtest):

If you don't know what a logarithm is you are gonna have a really tough time here. I think you need to go back and review what that is.

OpenStudy (anonymous):

\[a=b^x\] \[x=\log_ba\]

OpenStudy (anonymous):

we are going to introduce the logs both sides by writing them before the powers,is that okay

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[\log7^{t+1}=\log8^t\]

OpenStudy (anonymous):

look at @TuringTest rule,1st comment

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[(t+1)\log7=t \log8\]

OpenStudy (anonymous):

\[\frac{ t+1 }{ t }=\frac{ \log 8 }{ \log7}\]

OpenStudy (anonymous):

so did you get .388?

OpenStudy (anonymous):

OR\[7^t7=8^t\] \[(\frac{ 8 }{ 7 })^t=7\] \[t=\log_{\frac{ 8 }{ 7 }}7\]

OpenStudy (kingstone):

Sorry, I'm too lazy to show you my work, but.. the final answer is: \[t= \frac{ \log(7) }{ 3\log(2) - \log(7) }\]

OpenStudy (anonymous):

okay thanks for the help guys

OpenStudy (kingstone):

Just take my answer if you want the right answer..